# thanks: Brinson Prize Fellow

### v2

@@ Abstract
We present a measurement of the free-streaming length of dark matter (DM) and subhalo abundance around 28 quadruple image strong lenses using observations from JWST MIRI presented in Paper III of this series. We improve on previous inferences on DM properties from lensed quasars by simultaneously reconstructing extended lensed arcs with image positions and relative magnifications (flux ratios). Our forward modeling framework generates full populations of subhalos, line-of-sight halos, and globular clusters, uses an accurate model for subhalo tidal evolution, and accounts for free-streaming effects on halo abundance and concentration. Modeling lensed arcs leads to more-precise model-predicted flux ratios, breaking covariance between subhalo abundance and the free-streaming scale parameterized by the half-mode mass $m_{\rm{hm}}$. Assuming subhalo abundance predicted by the semi-analytic model galacticus ($N$-body simulations), we infer (Bayes factor of 10:1) $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ ($m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$), a 0.4 dex (0.3 dex) improvement relative to omitting lensed arcs. These bounds correspond to lower limits on thermal relic DM particle masses of $7.4$ and $8.4$ keV, respectively. Conversely, assuming DM is cold, we infer a projected mass in subhalos ($10^{6}<m/M_{\odot}<10^{10.7}$) of $1.6_{-1.1}^{+2.4}\times 10^{7}\ \mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ at $95\%$ confidence. This is consistent with galacticus predictions ($0.6\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$), but in mild tension with recent $N$-body simulations ($0.3\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$). Our results are among the strongest bounds on WDM, and the most precise measurement of subhalo abundance around strong lenses. Further improvements will follow from the large sample of lenses to be discovered by Euclid, Rubin, and Roman.
@@ II.3 Imaging data reconstruction
There is not a unique way to choose $\sigma_{\rm{noise}}$. We have experimented with defining this term in a way that connects directly to the uncertainties in the imaging data on a pixel level. For example, we can generate 10,000 background noise realizations (using $\sigma_{ii,(\rm{rms})}$ and $\sigma_{ii,\rm{(Poisson)}}$), render them on top of a best-fit lens model for each system, and set $\sigma_{ii,(\rm{noise})}$ equal to the standard deviations of the resulting distribution of log-likelihoods. However, strategies for defining $\sigma_{(\rm{noise})}$ in a way that connects to the background and Poison noise in the image tends to over-correct, in the sense that we discard more information from the imaging data that is necessary to make the problem computationally tractable. Instead, since the primary motivation for including $\sigma_{(\rm{noise})}$ is related to available CPU hours, $n_{\rm{cpu-hour}}$, we define $\sigma_{(\rm{noise})}$ in such a way that $\sigma_{(\rm{noise})}\rightarrow 0$ (meaning we do not down-weight constraints from lensed arcs) as $n_{\rm{cpu-hour}}\rightarrow\infty$. To compute $w_{\rm{img}}$, we generate $N\sim 200,000$ lens models per system, on average, and set $\sigma_{(\rm{noise})}$ such that $N/N_{\rm{eff}}=2000$. The threshold $N/N_{\rm{eff}}=2000$ represents a compromise between extracting more information from the lensed arcs, and having enough sampling of the parameter space to suppress shot-noise in the $w_{\rm{img}}$ distributions. We offer a complementary interpretation of the $\sigma_{(\rm{noise})}$ parameter: this parameter effectively bootstraps the uncertainties in the imaging data until the top $\sim 1\%$ of lens models have comparable log-likelihoods.
@@ III Datasets
$\Sigma_{\rm{sub}}\ \left[\rm{kpc^{-2}}\right]$ amplitude of the differential $\log_{10}\mathcal{U}\left(-2.2,0.2\right)$ NN-body predicts $\sim 0.05\ \rm{kpc^{-2}}$
subhalo mass function at infall SAM predicts $\sim 0.1\ \rm{kpc^{-2}}$
@@ IV.1.2 Subhalos
Equation 15 specifies the number density of halos with a mass $m$ at infall that eventually appear within a $30\ \rm{kpc}$ aperture of the host halo center (the region relevant for strong lensing). After generating this infall population, we use the tidal stripping framework presented by Du++25 to predict each subhalo’s density profile. The calculation begins with assigning infall redshifts to each subhalo using the probability distribution $p\left(z_{\rm{infall}}|m,m_{\rm{host}}\right)$, where the infall redshift depends on $m$ and the host-halo mass $m_{\rm{host}}$. From $z_{\rm{infall}}$ and $m$ we evaluate the concentration of the subhalo at infall $c=c\left(m,z_{\rm{infall}}\right)$. Next, given the host halo concentration $c_{\rm{host}}=c_{\rm{host}}\left(m_{\rm{host}},z_{\rm{d}}\right)$, which we calculate using the DiemerJoyce19 concentration–mass relation, we sample the bound mass fraction, $f_{\rm{bound}}\equiv m_{\rm{bound}}/m$, from the probability distribution $p\left(f_{\rm{bound}}|c,z_{\rm{infall}},c_{\rm{host}}\right)$. Du++25 derived empirical representations of this distribution by evolving subhalos in a growing host potential with galacticus. As discussed by Du++25, a given set of subhalo infall time, concentration, and host halo concentration specifies a distribution of $f_{\rm{bound}}$ due to differences in subhalo orbits and mass loss histories. Given $f_{\rm{bound}}$, we use tidal tracks [e.g. Errani++21, Du++24] to calculate the truncation radius $r_{\mathrm{t}}$, and a factor $f_{\rm{t}}$ that changes the normalization of the NFW profile $\rho_{\mathrm{s}}\rightarrow f_{\rm{t}}\rho_{\mathrm{s}}$. This model predicts that, at the redshift of the lens, most subhalos have lost over $90\%$ of the mass they had at infall. The model implemented in $\tt{pyHalo}$ shows excellent agreement with cosmological simulations of subhalo evolution performed with galacticus for $\Sigma_{\rm{sub}}\approx 0.1\ \rm{kpc^{-2}}$ [Du++25].
@@ IV.3 Sampling distributions and priors for dark matter hyper-parameters
We use a uniform sampling distribution on $\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}$ in the range $\left[4,10\right]$, and quote our constraints on this parameter with the same Bayesian prior. While CDM encompasses models with halos less massive than $10^{4}M_{\odot}$, and therefore $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ is not formally consistent with the theory, for $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ the abundance and density profiles of halos deviate by less than $10\%$ at $10^{6}\mathrm{M}_{\odot}$, the minimum halo mass rendered in our simulations. Therefore, we can identify regions of parameter space with $m_{\rm{hm}}<\mathcal{O}\left(10^{5}\mathrm{M}_{\odot}\right)$ as effectively indistinguishable from CDM.
@@
We draw $\Sigma_{\rm{sub}}/\rm{kpc^{-2}}$ from a log-uniform sampling distribution in $\log_{10}\Sigma_{\rm{sub}}\sim\left[-2.2,0.2\right]$, which spans a broad range around the theoretical predictions from the Symphony suite of $N$-body simulations. The value of $\Sigma_{\rm{sub}}$ that reproduces the galacticus bound mass function, given our tidal evolution model, and an additional suppression of $15\%$ due to the presence of the central elliptical galaxy [Gannon++25], is $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-2}}$ [Du++25]. The bound mass function predicted by Symphony is lower by a factor of 2 relative to the galacticus bound mass function [Gannon++25]. This implies $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ as the prediction from recent Symphony N-body simulations [Nadler++23]. In Section VI, we will present constraints on $m_{\rm{hm}}$ that use priors centered on these values with a factor of 2 uncertainty. We note that the analysis by Gannon++25 used the public dataset from the Symphony simulations, which makes use of the latest generation of halo finders [Mansfield++24]. The galacticus semi-analytic model was calibrated to the caterpillar suite of $N$-body simulations [Griffen++16].
Relative to previous analyses [Gilman++20, Keeley++24], we extend the sampling distribution on $\Sigma_{\rm{sub}}$ to significantly higher infall mass function amplitudes than these theoretical predictions to test the dependence of the results on the prior, and to demonstrate how imaging data helps break covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Finally, for each lens we sample the host halo mass from a log-normal prior $\log_{10}m_{\rm{host}}/M_{\odot}\sim\mathcal{N}\left(13.3,0.3\right)$, based on the typical host halo masses for strong lens systems [Lagattuta++10]. We marginalize over $m_{\rm{host}}$ for each lens before combining likelihoods, meaning the host halo mass is treated as a nuisance parameter in Equation 3.
@@ VI.2 Constraints on WDM
Throughout this section, we will quote the results of our analysis in terms of both the one-sided $95\%$ exclusion limit on $m_{\rm{hm}}$ obtained from the posterior, and the value of $m_{\rm{hm}}$ that corresponds to a Bayes factor (or relative likelihood) of 10:1 with respect to the peak of the posterior distribution. The motivation for quoting a Bayes factor is that this metric does not depend on the $m_{\rm{hm}}$ prior. On the other hand, for an inference that is consistent with CDM, we expect $m_{\rm{hm}}$ to be unconstrained from below, because the data are insensitive to $m_{\rm{hm}}\lesssim 10^{5}\mathrm{M}_{\odot}$. Confidence intervals and exclusion limits will therefore depend on the lower bound of the $m_{\rm{hm}}$ prior. While we will quote the one-sided exclusion limit, we emphasize that comparisons between difference constraints on $m_{\rm{hm}}$ in terms of a one-sided exclusion limit should be made using the same prior.
@@
From Figure 19 we see that $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ are correlated. The physical interpretation of this correlation is that adding additional more-massive subhalos to warmer (larger $m_{\rm{hm}}$) dark matter models can mimic the lensing signal of a colder dark matter model with fewer subhalos. Incorporating constraints from imaging data aids in breaking this covariance, strengthening the bound on $m_{\rm{hm}}$ by 0.7 dex, as summarized in the figure caption.
@@
Figure 19: Left:Posterior distribution of the hyper-parameters $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, the normalization of the differential subhalo mass function (Equation 15) and the half-mode mass (Equation 18), respectively. We marginalize over $\alpha$ and $\delta_{\rm{LOS}}$, the logarithmic slope of the subhalo mass function and the amplitude of the field halo mass function, respectively, as these parameters are unconstrained within the range of their priors. The black distribution uses only image positions and flux ratios, and the blue distribution shows the inference after incorporating imaging data. Right: Inference on $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ using flux ratios and lensed arcs, showing the effect of including a prior on $\Sigma_{\rm{sub}}$. The blue distribution is the same as the blue distribution in the left panel, and uses a log-uniform prior. The green distribution uses a prior $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\sim\mathcal{N}\left(-1.0,0.3\right)$ based on predictions from the semi-analytic model galacticus. The purple distribution shows the result when using a prior based on $N$-body simulations $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\sim\mathcal{N}\left(-1.3,0.3\right)$
NN-body simulations and semi-analytic models make theoretical predictions for the normalization of the subhalo mass function. However, $N$-body simulations and the semi-analytic model galacticus differ in their predictions by a factor of $\sim 2$. The $N$-body simulations presented by Nadler++23 predict $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ [Gannon++25], while the semi-analytic model galacticus predicts $\Sigma_{\rm{sub}}\sim 0.1\ \rm{kpc^{-2}}$ [Gannon++25, Du++25] (see discussion in Section IV.3).
@@
The right panel of Figure 19 shows the posterior distribution with informative priors on $\Sigma_{\rm{sub}}$ based on $N$-body simulations (purple) and semi-analytic model galacticus (green), as indicated in the figure caption. These priors break covariance between $\Sigma_{\rm{sub}}$ and the half-mode mass $m_{\rm{hm}}$. Assuming a prior $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.0,0.3\right)$, as predicted by galacticus, the bound on $m_{\rm{hm}}$ strengthens to $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$. This constraint is 0.4 dex stronger than the limit obtained without using lensed arcs, presented in a companion paper, Keeley++25. Assuming $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.3,0.3\right)$, as predicted by $N$-body simulations, we infer $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, an improvement of 0.3 dex relative to omitting imaging data. These upper limits on $m_{\rm{hm}}$ correspond to (see Equation 18) lower bounds on the mass of a thermal relic WDM particle of $7.4\ \rm{keV}$ and $8.4\ \rm{keV}$ for the galacticus and $N$-body priors, respectively. The corresponding one-sided $95\%$ exclusion limits are $m_{\rm{hm}}<10^{6.9}M_{\odot}$ and $m_{\rm{hm}}<10^{6.7}M_{\odot}$, respectively, which correspond to thermal relic particle constraints $m_{\rm{thermal}}>10.4\ \rm{keV}$ and $m_{\rm{thermal}}>11.9\ \rm{keV}$, respectively. As discussed in Section IV.2 (immediately following Equation 18) a slight modification to the conversion between $m_{\rm{hm}}$ and $m_{\rm{therm}}$ presented by Nadler++25cozmic gives bounds of $6.7$ keV and $7.6$ keV for the two Bayes factor limits.

### v3

@@ Abstract
We present a measurement of the free-streaming length of dark matter (DM) and subhalo abundance around 28 quadruple image strong lenses using observations from JWST MIRI presented in Paper III of this series. We improve on previous inferences on DM properties from lensed quasars by simultaneously reconstructing extended lensed arcs with image positions and relative magnifications (flux ratios). Our forward modeling framework generates full populations of subhalos, line-of-sight halos, and globular clusters, uses an accurate model for subhalo tidal evolution, and accounts for free-streaming effects on halo abundance and concentration. Modeling lensed arcs leads to more-precise model-predicted flux ratios, breaking covariance between subhalo abundance and the free-streaming scale parameterized by the half-mode mass $m_{\rm{hm}}$. Assuming subhalo abundance predicted by the semi-analytic model galacticus ($N$-body simulations), we infer (Bayes factor of 10:1) $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ ($m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$), a 0.4 dex improvement relative to omitting lensed arcs. These bounds correspond to lower limits on thermal relic DM particle masses of $7.4$ and $8.4$ keV, respectively. Conversely, assuming DM is cold, we infer a projected mass in subhalos ($10^{6}<m/M_{\odot}<10^{10.7}$) of $1.7_{-1.2}^{+2.6}\times 10^{7}\ \mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ at $95\%$ confidence. This is consistent with galacticus predictions ($0.9\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$), but in mild tension with recent $N$-body simulations ($0.6\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$). Our results are among the strongest bounds on WDM, and the most precise measurement of subhalo abundance around strong lenses. Further improvements will follow from the large sample of lenses to be discovered by Euclid, Rubin, and Roman.
@@ II.3 Imaging data reconstruction
There is not a unique way to choose $\sigma_{\rm{noise}}$. We have experimented with defining this term in a way that connects directly to the uncertainties in the imaging data on a pixel level. For example, we can generate 10,000 background noise realizations (using $\sigma_{ii,(\rm{rms})}$ and $\sigma_{ii,\rm{(Poisson)}}$), render them on top of a best-fit lens model for each system, and set $\sigma_{ii,(\rm{noise})}$ equal to the standard deviations of the resulting distribution of log-likelihoods. However, strategies for defining $\sigma_{(\rm{noise})}$ in a way that connects to the background and Poison noise in the image tends to over-correct, in the sense that we discard more information from the imaging data that is necessary to make the problem computationally tractable. Instead, since the primary motivation for including $\sigma_{(\rm{noise})}$ is related to available CPU hours, $n_{\rm{cpu-hour}}$, we define $\sigma_{(\rm{noise})}$ in such a way that $\sigma_{(\rm{noise})}\rightarrow 0$ (meaning we do not down-weight constraints from lensed arcs) as $n_{\rm{cpu-hour}}\rightarrow\infty$. To compute $w_{\rm{img}}$, we generate $N\sim 200,000$ lens models per system, on average, and set $\sigma_{(\rm{noise})}$ such that $N/N_{\rm{eff}}=2000$. The threshold $N/N_{\rm{eff}}=2000$ represents a compromise between extracting more information from the lensed arcs, and having enough sampling of the parameter space to suppress shot-noise in the $w_{\rm{img}}$ distributions. This approach effectively bootstraps uncertainties in the imaging data through $\sigma_{\rm{noise}}$ until the top $\sim 1\%$ of lens models have comparable log-likelihoods.
@@ III Datasets
$\Sigma_{\rm{sub}}\ \left[\rm{kpc^{-2}}\right]$ amplitude of the differential $\log_{10}\mathcal{U}\left(-2.2,0.2\right)$ NN-body predicts $\sim 0.1\ \rm{kpc^{-2}}$
subhalo mass function at infall SAM predicts $\sim 0.15\ \rm{kpc^{-2}}$
@@ IV.1.2 Subhalos
Equation 15 specifies the number density of halos with a mass $m$ at infall that eventually appear within a $30\ \rm{kpc}$ aperture of the host halo center (the region relevant for strong lensing). After generating this infall population, we use the tidal stripping framework presented by Du++25 to predict each subhalo’s density profile. The calculation begins with assigning infall redshifts to each subhalo using the probability distribution $p\left(z_{\rm{infall}}|m,m_{\rm{host}}\right)$, where the infall redshift depends on $m$ and the host-halo mass $m_{\rm{host}}$. From $z_{\rm{infall}}$ and $m$ we evaluate the concentration of the subhalo at infall $c=c\left(m,z_{\rm{infall}}\right)$. Next, given the host halo concentration $c_{\rm{host}}=c_{\rm{host}}\left(m_{\rm{host}},z_{\rm{d}}\right)$, which we calculate using the DiemerJoyce19 concentration–mass relation, we sample the bound mass fraction, $f_{\rm{bound}}\equiv m_{\rm{bound}}/m$, from the probability distribution $p\left(f_{\rm{bound}}|c,z_{\rm{infall}},c_{\rm{host}}\right)$. Du++25 derived empirical representations of this distribution by evolving subhalos in a growing host potential with galacticus. As discussed by Du++25, a given set of subhalo infall time, concentration, and host halo concentration specifies a distribution of $f_{\rm{bound}}$ due to differences in subhalo orbits and mass loss histories. Given $f_{\rm{bound}}$, we use tidal tracks [e.g. Errani++21, Du++24] to calculate the truncation radius $r_{\mathrm{t}}$, and a factor $f_{\rm{t}}$ that changes the normalization of the NFW profile $\rho_{\mathrm{s}}\rightarrow f_{\rm{t}}\rho_{\mathrm{s}}$. This model predicts that, at the redshift of the lens, most subhalos have lost over $90\%$ of the mass they had at infall. The model implemented in $\tt{pyHalo}$ shows excellent agreement with cosmological dark matter only simulations of subhalo evolution performed with galacticus for $\Sigma_{\rm{sub}}\approx 0.17\ \rm{kpc^{-2}}$ [Du++25]. In Section IV.3, we discuss theoretical priors on $\Sigma_{\rm{sub}}$ based on both $\tt{galacticus}$ and the Symphony $N$-body simulations.
@@ IV.3 Sampling distributions and priors for dark matter hyper-parameters
We use a uniform sampling distribution on $\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}$ in the range $\left[4,10\right]$, and quote our constraints on this parameter with the same Bayesian prior. While CDM encompasses models with halos less massive than $10^{4}M_{\odot}$, and therefore $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ is not formally consistent with the theory, for $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ the abundance and density profiles of halos deviate by less than $10\%$ at $10^{6}\mathrm{M}_{\odot}$, the minimum halo mass rendered in our simulations. Therefore, we can identify regions of parameter space with $m_{\rm{hm}}\lesssim\mathcal{O}\left(10^{5}\mathrm{M}_{\odot}\right)$ as effectively indistinguishable from CDM.
@@
We draw $\Sigma_{\rm{sub}}/\rm{kpc^{-2}}$ from a log-uniform sampling distribution in $\log_{10}\Sigma_{\rm{sub}}\sim\left[-2.2,0.2\right]$, which spans a broad range around the theoretical predictions from the Symphony suite of $N$-body simulations. Relative to previous analyses [Gilman++20, Keeley++24], we extend the sampling distribution on $\Sigma_{\rm{sub}}$ to significantly higher infall mass function amplitudes to test the dependence of the results on the prior, and to demonstrate how imaging data helps break covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$.
In Section VI, we will incorporate theoretically-motivated priors on $\Sigma_{\rm{sub}}$ from both galacticus and the Symphony $N$-body simulations [Nadler++23] when quoting constraints on $m_{\rm{hm}}$. We will also compare measurements of subhalo abundance in CDM with the theoretical predictions. We determine the theoretical predictions for $\Sigma_{\rm{sub}}$ by examining the bound mass function in dark-matter-only simulations in a 10 kpc annulus around the center of the hosts [Gannon++25]. Specifically, we determine the prior on $\Sigma_{\rm{sub}}$, the amplitude of the mass function at infall, that results in the bound mass function predicted by either Symphony or $\tt{galacticus}$}, given our tidal evolution model [Du++25]. From our tidal evolution model, in CDM, for a group scale halo of 1013 M at redshift 0.5, the bound mass function is related to the infall mass function by a constant factor, which we refer to as the mean bound mass fraction, $\bar{f}_{\rm{bound}}=0.05$ [Du++25]. When determining priors on $\Sigma_{\rm{sub}}$, we also take into account the additional $15\%$ suppression of the bound mass function amplitude caused by tidal stripping effects from a central elliptical galaxy [Gannon++25].
The galacticus dark-matter-only prediction for the bound mass function amplitude corresponds to $\bar{f}_{\rm{bound}}\times\Sigma_{\rm{sub}}=0.009\ \rm{kpc^{-2}}$. Including an additional $15\%$ suppression due to tidal stripping by the central galaxy gives $0.0075\ \rm{kpc^{-2}}$. This gives $\Sigma_{\rm{sub}}=0.15\ \rm{kpc^{-2}}$ as the prediction corresponding to $\tt{galacticus}$. We note that because we use galacticus to calibrate the tidal evolution model, by construction we also match the predicted infall mass function amplitude. This is not the case for the $N$-body simulations, for which we only have knowledge of the bound mass function amplitude. As discussed by Gannon++25, the Symphony $N$-body simulations, analyzed using the latest generation of halo finders [Mansfield++24], predict a bound mass function amplitude suppressed by $\sim 50\%$ relative to galacticus in the inner regions of the host, corresponding to $\bar{f}_{\rm{bound}}\times\Sigma_{\rm{sub}}=0.006\ \rm{kpc^{-2}}$. Including the additional $15\%$ suppression from the central galaxy, we find $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-2}}$ gives the amplitude of the bound mass function predicted by $N$-body simulations. Based on these considerations, we use theoretically-motivated priors based on galacticus $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-0.8,0.2\right)$ and Symphony $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.0,0.2\right)$. The width of these priors is motivated by the 0.2 dex difference between the predicted mean $\Sigma_{\rm{sub}}$ values, which we take as a representative level of systematic uncertainty between different tools for examining non-linear structure formation.
Finally, for each lens we sample the host halo mass from a log-normal prior $\log_{10}m_{\rm{host}}/M_{\odot}\sim\mathcal{N}\left(13.3,0.3\right)$, based on the typical host halo masses for strong lens systems [Lagattuta++10]. We marginalize over $m_{\rm{host}}$ for each lens before multiplying likelihoods, meaning the host halo mass is treated as a nuisance parameter in Equation 3.
@@ VI.2 Constraints on WDM
Throughout this section, we will quote the results of our analysis in terms of both the one-sided $95\%$ exclusion limit on $m_{\rm{hm}}$ obtained from the posterior, and the value of $m_{\rm{hm}}$ that corresponds to a Bayes factor (or relative likelihood) of 10:1 with respect to the peak of the posterior distribution. The motivation for quoting a Bayes factor is that this metric does not depend on the $m_{\rm{hm}}$ prior. On the other hand, for an inference that is consistent with CDM, we expect $m_{\rm{hm}}$ to be unconstrained from below, because the data are insensitive to $m_{\rm{hm}}\lesssim 10^{5}\mathrm{M}_{\odot}$. Confidence intervals and exclusion limits will therefore depend on the lower bound of the $m_{\rm{hm}}$ prior. Any future comparison with our bounds on $m_{\rm{hm}}$ in terms of a one-sided exclusion limit should be made using the same prior.
@@
From Figure 19 we see that $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ are correlated. The physical interpretation of this correlation is that adding additional more-massive subhalos to warmer (larger $m_{\rm{hm}}$) dark matter models can mimic the lensing signal of a colder dark matter model with fewer subhalos. Incorporating constraints from imaging data aids in breaking this covariance, strengthening the bound on $m_{\rm{hm}}$ by 0.6 dex, as summarized in the figure caption.
@@
Figure 19: Left:Posterior distribution of the hyper-parameters $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, the normalization of the differential subhalo mass function (Equation 15) and the half-mode mass (Equation 18), respectively. We marginalize over $\alpha$ and $\delta_{\rm{LOS}}$, the logarithmic slope of the subhalo mass function and the amplitude of the field halo mass function, respectively, as these parameters are unconstrained within the range of their priors. The black distribution uses only image positions and flux ratios, and the blue distribution shows the inference after incorporating imaging data. Right: Inference on $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ using flux ratios and lensed arcs, showing the effect of including a prior on $\Sigma_{\rm{sub}}$. The blue distribution is the same as the blue distribution in the left panel, and uses a log-uniform prior. The green distribution uses a prior $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\sim\mathcal{N}\left(-0.8,0.2\right)$ based on predictions from the semi-analytic model galacticus. The purple distribution shows the result when using a prior based on $N$-body simulations $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\sim\mathcal{N}\left(-1.0,0.2\right)$, as discussed in Section IV.3
@@
The right panel of Figure 19 shows the posterior distribution with informative priors on $\Sigma_{\rm{sub}}$ based on $N$-body simulations (purple) and semi-analytic model galacticus (green), as indicated in the figure caption. These priors break covariance between $\Sigma_{\rm{sub}}$ and the half-mode mass $m_{\rm{hm}}$. As discussed in Section IV.3, we derive these priors by determining the value of $\Sigma_{\rm{sub}}$ (the infall mass function amplitude) that produces the bound mass function amplitude predicted by $\tt{galacticus}$} and the Symphony $N$-body simulations [Nadler++23]. For a prior $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-0.8,0.2\right)$, as predicted by galacticus, the bound on $m_{\rm{hm}}$ strengthens to $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$. This constraint is 0.5 dex stronger than the limit obtained without using lensed arcs, presented in a companion paper, Keeley++25. Assuming $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.0,0.2\right)$, as predicted by $N$-body simulations, we infer $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, an improvement of 0.4 dex relative to omitting imaging data. These upper limits on $m_{\rm{hm}}$ correspond to (see Equation 18) lower bounds on the mass of a thermal relic WDM particle of $7.4\ \rm{keV}$ and $8.4\ \rm{keV}$ for the galacticus and $N$-body priors, respectively. The corresponding one-sided $95\%$ exclusion limits are $m_{\rm{hm}}<10^{6.8}M_{\odot}$ and $m_{\rm{hm}}<10^{6.7}M_{\odot}$, respectively, which correspond to thermal relic particle constraints $m_{\rm{thermal}}>11.1\ \rm{keV}$ and $m_{\rm{thermal}}>11.9\ \rm{keV}$, respectively. As discussed in Section IV.2 (immediately following Equation 18) a slight modification to the conversion between $m_{\rm{hm}}$ and $m_{\rm{therm}}$ presented by Nadler++25cozmic gives bounds of $6.7$ keV and $7.6$ keV for the two Bayes factor limits.